OpenStudy (anonymous):

how to write an equation in slope-intercept form of the line that passes through the given points? (3,5) and (0,4) #HELP

OpenStudy (anonymous):

first find the slope. then use the equation y = mx +b with one point to find the b and then u have your answer

OpenStudy (amistre64):

given a point (Px,Py) and a slope m: y=mx-mPx+Py

OpenStudy (amistre64):

thought one of those was a slope :)

OpenStudy (amistre64):

subtract one point from the other and stack y/x for slope

OpenStudy (anonymous):

and how do you do that?

OpenStudy (anonymous):

could you solve this one and show me the work, so i can follow it? :)

OpenStudy (amistre64):

(3,5) -(0,4) ------ x,y

OpenStudy (anonymous):

slope = (5-4)/(3-0)

OpenStudy (anonymous):

so that is the final answer? :)

OpenStudy (amistre64):

try it out and see :)

OpenStudy (anonymous):

if i knew how to.

OpenStudy (amistre64):

when x=0, does y=4? when x=3 does y=5?